# The Geometry of Queries: Query-Based Innovations in Retrieval-Augmented Generation

### v1

@@ 5.4 Case Study
6 Discussion
This paper introduces QB-RAG, a novel approach for enhancing the retrieval phase of RAG systems. QB-RAG addresses the inherent semantic misalignment between user queries and knowledge bases by pre-generating a comprehensive set of potential questions directly from the content. We then leverage efficient vector search to map incoming online queries to these pre-computed questions, facilitating a more accurate and aligned retrieval process. While QB-RAG requires a significant offline computational investment for question generation, this trade-off is strategically advantageous. Unlike conventional methods that rely on online LLM calls for query rewriting or enhancement, QB-RAG shifts this computational burden offline. This distinction is crucial for minimizing latency in real-time applications, as online LLM invocations can significantly impact the user experience. Furthermore, QB-RAG benefits from the efficiency of highly optimized retrieval algorithms, even when handling large question sets, ensuring rapid content retrieval for online queries. We have provided a theoretical motivations for QB-RAG and we have also suggested potential avenues for further improving QB-RAG, such as expanding the question set for greater coverage or leveraging matrix completion techniques to infer a complete answerability matrix.
To assess QB-RAG’s effectiveness, we designed two distinct test sets: the Rephrase set, where test questions are semantically similar to those in our pre-generated question base, and the Out-of-Distribution set, where questions are intentionally crafted to be dissimilar. These test sets represent cases designed to evaluate QB-RAG’s robustness under varying conditions, including a potential distribution shift. Our empirical results demonstrate that QB-RAG consistently outperforms benchmark methods across both test sets. While QB-RAG significantly improves exact recovery and relevance of retrieved documents, its impact extends beyond these initial retrieval metrics. Downstream answer generation using QB-RAG-retrieved content scores higher on key quality metrics, including faithfulness, relevancy, and adherence to pre-defined guidelines. Crucially, answers generated using QB-RAG are more frequently grounded on our trusted source of content, which is key to deliver reliable, up to date information to patient queries.
Our sensitivity analysis shows that effectiveness of our approach is highly tied to coverage of extracted questions. Beyond increasing the raw number of questions, we also suggest favoring a diverse set of questions during the generation phase. This can be done potentially with prompt engineering, supervised fine-tuning or RLHF frameworks. Metrics like BERTScore (Zhang et al., 2019) and Word Mover’s Distance (Kusner et al., 2015) can be used to evaluate the diversity of generated questions from the RAG knowledge base. Finally, an LLM could also be potentially used as an auto-evaluator for measuring diversity of generated questions. We note that retrieval frameworks are highly optimized even for large question bases, thus, we don’t expect quantity of generated questions as a bottleneck for many real-world applications.
Despite its novelty and effectiveness, our work has several limitations that require further study. First, our evaluation focused on a curated healthcare content base consisting of well-crafted, concise pieces. Assessing QB-RAG’s performance on larger, more diverse datasets, including those from other domains and containing less structured or longer-form content, remains an important area for future research.
Another limitation stems from our reliance on LLM-based evaluation metrics in the current study. While LLMs have shown promise in assessing answer quality according to predefined guidelines, the real-world deployment of QB-RAG in healthcare settings necessitates thorough expert review. Direct patient interactions demand the highest level of scrutiny to ensure the accuracy, safety, and reliability of generated information. Our automated evaluation serves as a valuable first step, but expert validation remains paramount.
Finally, it’s crucial to acknowledge a broader challenge inherent to all RAG systems, particularly in fast-changing domains like healthcare: the quality of any RAG system’s output is fundamentally dependent on the quality and recency of its underlying knowledge base. In healthcare, where medical consensus and best practices are constantly evolving, content needs to be regularly reviewed and updated to reflect the latest advancements. This underscores the importance of developing robust mechanisms for ongoing knowledge base maintenance and ensuring alignment with current clinical guidelines, regardless of the specific retrieval methods employed.
Despite these limitations, our work demonstrates the potential of QB-RAG for enhancing retrieval accuracy and answer groundedness in RAG systems, paving the way for more reliable and trustworthy LLM applications in healthcare. Beyond direct patient interaction, QB-RAG could also be valuable for augmenting coaching and provider resources within digital health platforms. For instance, QB-RAG could be used to generate draft answers that healthcare professionals can then review and refine, ensuring both efficiency and accuracy. Ultimately, real-world deployment of QB-RAG should be carefully considered, taking into account specific use cases, potential risks, and the need for appropriate expert oversight.

### v2

@@ Appendix G Case Study
Appendix H Prompts
H.1 Prompt for QB-RAG Question Generation
***SYSTEM:***
You are a Teacher / Professor. Your task is to setup {num_questions} questions for an upcoming quiz/examination. The questions should be both diverse and exhaustive in nature across the document. Restrict the questions to the context information provided.
***INSTRUCTIONS:***
You are presented with a text authored by healthcare professionals, offering advice and strategies for managing conditions such as diabetes and high blood pressure. Your task is to formulate relevant questions that the text is written to address. Closely follow the example questions for style and structure when formulating your own question for the provided text. Your generated questions should be in first person with details, but only at a high school reading level. Your questions should be answerable from the text, but do not copy the text verbatim. MAKE SURE to generate at least {num_questions} questions. Format the generated questions separated by comma in the following JSON format with "questions" as its key: {{"questions": ["...","...","...","..."]}}
***EXAMPLE OUTPUTS:***
Sample Text: {example content}
Generated Questions: {example JSON with list of num_question questions}
Given the context information and no prior knowledge, generate the relevant questions.
Text:
{cc_text}
Generated Questions:
Listing 1: Base prompt for question generation
H.2 Prompt for Answerability Auto-Evaluator
***SYSTEM***
You are a health coach providing support for members living with diabetes. You have some basic healthcare and nutrition knowledge.
***INSTRUCTIONS***
Given a pair of user query and a paragraph of content, determine if the content contains relevant information to infer an answer to the query. Think step by step. First provide an explanation, then generate a "Yes" or "No" label. Put the results in a Python dictionary format with keys "Explanation" and "Source relevant".
***EXAMPLES***
For example, given the following query and content as inputs: {positive and negative examples, with explanations}
Provide the output for the following query and content:
Question: {question}
Content: {content}
Listing 2: Prompt for Answerability Auto-Evaluator
H.3 Prompt for Answer Generation
Use only the provided pieces of context to answer the question at the end. Think step-by-step and then answer. Respond in 3 to 6 short sentences.
Do not try to make up an answer:
- If the context does not contain enough relevant information to determine an answer to the query, say "I cannot determine the answer to that."
- If the context is empty, just say "I do not know the answer to that."
Answer in an empathetic and positive tone. Do not use phrases such as "According to the context", and directly answer the question.
Contexts: {contexts}
Question: {question}
Answer:
Listing 3: Base prompt for answer generation